a poaster is 25 cm taller than it is wide. It is mounted on a piece of cardboard so that there is a 5 cm boarder on all sides. If the area of the boarder alone is 1350 cm2, what are the dimensions of the poster?
Answer by mananth(16949)

a poaster is 25 cm taller than it is wide. It is mounted on a piece of cardboard so that there is a 5 cm boarder on all sides. If the area of the boarder alone is 1350 cm2, what are the dimensions of the poster?
...
width = w
length = w+25
..
width including border = w+10
length including border= w+25+10=w+35
...
(w+10)(w+35)-w(w+25)=1350
w^2+45w+350 -w^2-25w=1350
20w+350=1350
20w=1350-350
20w=1000
w=50 cm the width of poster
length = w+25=75 cm
...
m.ananth@hotmail.ca